# Close-Talk, and Pseudo-Label Based Far-Field, Speech Enhancement

### v1

@@ VI-B Synchronization of Close-talk and Far-field Mixtures
where $j$ denotes the imaginary unit, $d$ a hypothesized frame delay, and $\mathcal{R}(\cdot)$ extracts the real component. We then enumerate a set of hypothesized frame delays and find a delay that can produce the largest summation of the GCC-PHAT coefficients at all the far-field microphones and T-F units:
@@
VI-C Training Setup
For CTSEnet, it is trained based on all the $7,138\times 6$ monaural simulated mixtures. Following [3], we apply an SNR augmentation technique to the simulated training mixtures of CHiME-4. That is, during training, we optionally modify the SNR of each simulated mixture, on the fly, by $u$ dB, with $u$ uniformly sampled from the range $[-10,+15]$ dB. No other data augmentation is used.
For ctPuLSEnet, we use all the $7,138$ simulated and $1,600$ real utterances for training. For monaural ctPuLSEnet, we train it on all the $(7,138+1,600)\times 6$ monaural mixtures. For $2$-channel ctPuLSEnet, at each training step we sample two microphones from the front five microphones as input, and ctPuLSEnet is trained to predict the target speech at the first of the two selected microphones. For $6$-channel ctPuLSEnet, it stacks all the six microphones in a fixed order as input to predict the target speech at the fifth microphone. We apply the same SNR augmentation used in CTSEnet to the simulated training mixtures when training ctPuLSEnet. We always train ctPuLSEnet on the combination of the simulated and real mixtures using the co-learning algorithm introduced in Section V-D, as there are only $1,600$ real mixtures (which amount to only $\sim$2.72.72.72.7 hours) in the training set of CHiME-4
@@ VI-C Training Setup
VI-D Miscellaneous System Configurations
@@
TF-GridNet [38] is employed as the DNN architecture. It has shown strong separation and enhancement performance recently in serveral representitive supervised speech separation benchmarks. Following the symbols defined in Table I of [38], we set its hyper-parameters to $D=128$, $B=4$, $I=1$, $J=1$, $H=200$, $L=4$ and $E=4$ (please do not confuse the symbols with the ones defined in this paper). The model has $\sim$5.45.45.45.4 million trainable parameters. This configuration of TF-GridNet is used for both CTSEnet and ctPuLSEnet.
@@ VI-E ASR Model
The ASR model is pre-trained in a multi-conditional way on the official CHiME-4 simulated and real mixtures plus the clean speech signals in WSJ00 by using the script proposed in [44] [^3] https://github.com/espnet/espnet/blob/master/egs2/chime4/asr1/conf/tuning/train_asr_transformer_wavlm_lr1e-3_specaug_accum1_preenc128_warmup20k.yaml, which is available in the ESPnet toolkit. It is an encoder-decoder- and transformer-based system, trained on WavLM features [45] and using a transformer-based language model for decoding. From the results reported in [44] and [46], this model is the current best ASR model on CHiME-4.
We have successfully reproduced the ASR system in [44]. The mixture ASR results (shown later in row 00 of Table III) are very close to the ones reported in row $7$ of Table $1$ of [44].
VI-F Evaluation Metrics
@@
For simulated mixtures, where the clean speech is available, our evaluation metrics include short-time objective intelligibility (STOI) [47], wide-band perceptual evaluation of speech quality (WB-PESQ) [48], signal-to-distortion ratio (SDR) [49], and scale-invariant SDR (SI-SDR) [50]. They are designed to evaluate the intelligibility, quality, and accuracy of the magnitude and phase of enhanced speech. They are widely-adopted metrics in speech enhancement.
VI-G Baselines
@@
VI-H Tricks to Improve ASR Performance
@@ VII Evaluation Results
VII-A Results of Close-Talk Speech Enhancement
Table II presents the results of monaural CTSEnet.

### v2

@@ VI.2 Synchronization of Close-talk and Far-field Mixtures
where $j$ denotes the imaginary unit, $d\in{\mathbb{Z}}$ a hypothesized frame delay, and $\mathcal{R}(\cdot)$ extracts the real component. We then enumerate a set of hypothesized frame delays and find a delay that can produce the largest summation of the GCC-PHAT coefficients at all the far-field microphones and T-F units:
@@
VI.3 Training Setup
For CTSEnet, it is trained based on all the $7,138\times 6$ monaural simulated mixtures. Following [36], we apply an SNR augmentation technique to the simulated training mixtures of CHiME-4. That is, during training, we optionally modify the SNR of each simulated mixture, on the fly, by $u$ dB, with $u$ uniformly sampled from the range [10,+15][-10,+15] dB. No other data augmentation is used.
For ctPuLSEnet, we use all the $7,138$ simulated and $1,600$ real utterances for training. For monaural ctPuLSEnet, we train it on all the $(7,138+1,600)\times 6$ monaural mixtures. For $2$-channel ctPuLSEnet, at each training step we sample two microphones from the front five microphones as input, and ctPuLSEnet is trained to predict the target speech at the first of the two selected microphones. For $6$-channel ctPuLSEnet, it stacks all the six microphones in a fixed order as input to predict the target speech at the fifth microphone. We apply the same SNR augmentation used in CTSEnet to the simulated training mixtures when training ctPuLSEnet. We always train ctPuLSEnet on the combination of the simulated and real mixtures using the co-learning algorithm introduced in Section V.4, as there are only $1,600$ real mixtures (which amount to only $\sim$$2.7$ hours) in the training set of CHiME-4
@@ VI.3 Training Setup
VI.4 Miscellaneous System Configurations
@@
TF-GridNet [37] is employed as the DNN architecture. It has shown strong separation and enhancement performance recently in serveral representitive supervised speech separation benchmarks. Following the symbols defined in Table I of [37], we set its hyper-parameters to $D=128$, $B=4$, $I=1$, $J=1$, $H=200$, $L=4$ and $E=4$ (please do not confuse the symbols with the ones defined in this paper). The model has $\sim$$5.4$ million trainable parameters. This configuration of TF-GridNet is used for both CTSEnet and ctPuLSEnet.
@@ VI.5 ASR Model
The ASR model is pre-trained in a multi-conditional way on the official CHiME-4 simulated and real mixtures plus the clean speech signals in WSJ0 by using the script proposed in [5] [^1] https://github.com/espnet/espnet/blob/master/egs2/chime4/asr1/conf/tuning/train_asr_transformer_wavlm_lr1e-3_specaug_accum1_preenc128_warmup20k.yaml, which is available in the ESPnet toolkit. It is an encoder-decoder- and transformer-based system, trained on WavLM features [8] and using a transformer-based language model for decoding. From the results reported in [5] and [22], this model is the current best ASR model on CHiME-4.
We have successfully reproduced the ASR system in [5]. The mixture ASR results (shown later in row 0 of Table 3) are very close to the ones reported in row 77 of Table 11 of [5].
VI.6 Evaluation Metrics
@@
For simulated mixtures, where the clean speech is available, our evaluation metrics include short-time objective intelligibility (STOI) [13], wide-band perceptual evaluation of speech quality (wPESQ) [26], signal-to-distortion ratio (SDR) [19], and scale-invariant SDR (SISDR) [32]. They are designed to evaluate the intelligibility, quality, and accuracy of the magnitude and phase of enhanced speech. They are widely-adopted metrics in speech enhancement.
VI.7 Baselines
@@
VI.8 Tricks to Improve ASR Performance
@@ VII Evaluation Results
VII.1 Results of Close-Talk Speech Enhancement
Table 2 presents the results of monaural CTSEnet.